Mulligan Manufacturing Company uses a job order cost system with overhead applied to products at a rate of 150 percent of direct labor cost. Required: Treating each case independently, selected from the manufacturing data given below, find the missing amounts. You should do them in the order listed. (Hint: For the manufacturing costs in Case 3, first solve for conversion costs and then determine how much of that is direct labor and how much is manufacturing overhead.) (Do not round your intermediate calculations. Round your final answers to the nearest whole dollar. Enter all amounts as positive values.)

Case 1 Case 2 Case 3 10,500 Direct material used Direct labor Manufacturing overhead applied Total current manufacturing costs Beginning work in process inventory Ending work in process inventory Cost of goods manufactured Beginning finished goods inventory Ending finished goods inventory Cost of goods sold 18,000 18,000 13,000 27,500 31,800 7,500 8,500 9,600 50,000 11,200 4,400 12,001 3,600 7,500 6,700 41,000 35,000
Answer #1

Calculate following

Case 1 Case 2 Case 3
Direct material used 18000 5833 10500
Direct labor 18000 13000/1.5 = 8667 8520
Manufacturing overhead applied 18000*150% = 27000 13000 21300*1.5/2.5 = 12780
Total Current manufacturing cost 63000 27500 31800
Beginning work in process inventory 8500 32100 7500
Ending work in process inventory 4400 9600 27299
Cost of goods manufactured 67100 50000 12001
Beginning finished goods inventory 3600 11200 35699
Ending finished goods inventory 7500 26200 6700
Cost of goods sold 63200 35000 41000